Three-fourths of a number is 6. Find the number
Answer by Maths68(1474)

Let x be the number
3/4 of x = 6
Means
(3/4)(x)= 6
3x/4=6
Multiply by 4 both sides
(4)(3x/4)=(6)(4)
Cancel out 4 from Right hand side and multiply 4 on left hand side
3x=24
Divide by 3 both sides
3x/3=24/3
Cancel out 3 from Right hand side
x=8
Number is 8